# Egor Cherepanov1,2Nikita Kachaev1Alexey K. Kovalev1,2Aleksandr I. Panov1,2 1AIRI, Moscow, Russia  2MIPT, Dolgoprudny, Russia

### v1

@@ Appendix A MIKASA-Robo Implementation Details
An example of running the environment from the MIKASA-Robo benchmark is shown in 1. For ease of debugging, we also added various wrappers (found in mikasa_robo/utils/wrappers/) that display useful information about the episode on the video (2). Thus, RenderStepInfoWrapper() displays the current step in the environment; DebugRewardWrapper() displays information about the full reward at the current step in the environment; DebugRewardWrapper() displays information about each component that generates the reward function at the current step. In addition, we also added task-specific wrappers for each environment. For example, RememberColorInfoWrapper() displays the target color of the cube in the RememberColor-v0 task, and ShellGameRenderCupInfoWrapper() displays which mug the ball is actually under in the ShellGame-v0 task.
@@
1import mikasa_robo
2from mikasa_robo.utils.wrappers import *
3
4num_envs, seed = 512, 123
5
6# Create the environment via gym.make()
7# obs_mode="rgb" for modes "RGB", "RGB+joint", "RGB+oracle" etc.
8# obs_mode="state" for mode "state"
9env = gym.make("RememberColor9-v0", num_envs=num_envs,
10 obs_mode="rgb", render_mode="all")
11
12env = StateOnlyTensorToDictWrapper(env) # [always] gen. obs keys
13
14obs, _ = env.reset(seed)
15for i in tqdm(range(89)):
16 action = torch.from_numpy(env.action_space.sample())
17 obs, reward, terminated, truncated, info = env.step(action)
18env.close()
@@
1import mikasa_robo
2from mikasa_robo.utils.wrappers import *
3from mani_skill.utils.wrappers import RecordEpisode
4
5num_envs, seed = 512, 123
6env = gym.make("RememberColor9-v0", num_envs=num_envs,
7 obs_mode="rgb", render_mode="all")
8
9env = StateOnlyTensorToDictWrapper(env) # [always] gen. obs keys
10env = RememberColorInfoWrapper(env) # [debug] show task info
11env = RenderStepInfoWrapper(env) # [debug] show env step
12env = RenderRewardInfoWrapper(env) # [debug] show total reward
13env = DebugRewardWrapper(env) # [debug] show reward info
14env = RecordEpisode(env, "./videos/demo_remember-color-9")
15
16obs, _ = env.reset(seed)
17for i in tqdm(range(89)):
18 action = torch.from_numpy(env.action_space.sample())
19 obs, reward, terminated, truncated, info = env.step(action)
20env.close()
21
22Video("./videos/demo_remember-color-9/0.mp4", embed=True)
Appendix B MIKASA-Base Implementation Details
@@
1import mikasa_base
2import gymnasium as gym
3
4# use pre-defined env
5# env_id = "MemoryLengthEasy-v0"
6# env_kwargs = None
7
8# create env using custom parameters
9env_id = "MemoryLength-v0"
10env_kwargs = {"memory_length": 10, "num_bits": 1}
11seed = 123
12
13env = gym.make(env_id, env_kwargs)
14
15obs, _ = env.reset(seed=seed)
16
17for i in range(11):
18 action = env.action_space.sample()
19 next_obs, reward, terminations, truncations, infos = env.step(action)
20env.close()
Appendix C Memory Mechanisms in RL
@@
There are several approaches to incorporating memory into RL, including recurrent neural networks (RNNs) (Rumelhart et al., 1986; Hochreiter & Schmidhuber, 1997; Chung et al., 2014) which uses hidden states to store information from previous steps (Wierstra et al., 2010; Hausknecht & Stone, 2015), state-space models (SSMs) (Gu et al., 2021; Smith et al., 2023; Gu & Dao, 2023) which uses system state to store historical information (Hafner et al., 2019; Samsami et al., 2024), transformers (Vaswani et al., 2017) which uses attention mechanism to capture sequential dependencies inside the context window (Parisotto et al., 2020; Lampinen et al., 2021; Ni et al., 2023), graph neural networks (GNNs) (Zhou et al., 2020) which uses graphs to store information Zhu et al. (2023); Kang et al. (2024) etc. Popular agents with memory mechanisms are summarized in Table 2.
Appendix D Classic baselines performance on the MIKASA-Robo benchmark
@@
Appendix E MIKASA-Robo Detailed Tasks Description
@@ E.12 ChainOfColors-v0
Appendix F MIKASA-Base Benchmark Tasks Description
@@ F.1 Memory Cards
The Memory Cards environment (Esslinger et al., 2022) is a memory game environment with 5 randomly shuffled pairs of hidden cards. At each step, the agent sees one revealed card and must find its matching pair. A correct guess removes both cards; otherwise, the card is hidden again, and a new one is revealed. The game continues until all pairs are removed.
F.2 Numpad
@@
The Numpad environment (Humplik et al., 2019) consists of an $N\times N$ grid of tiles. The agent controls a ball that rolls between tiles. At the beginning of an episode, a random sequence of $n$ neighboring tiles (excluding diagonals) is selected, and the agent must follow this sequence in the correct order. The environment is structured so that pressing the correct tile lights it up, while pressing an incorrect tile resets progress. A reward of +1 is given for the first press of each correct tile after a reset. The episode ends after a fixed number of steps. To succeed, the agent must memorize the sequence and navigate it correctly without mistakes. The ability to “jump” over tiles is not available.
F.3 BSuite MemoryLength
@@
The MemoryLength environment (Osband et al., 2020) represents a sequence of observations, where at each step, the observation takes a value of either +1 or -1. The environment is structured so that a reward is given only at the final step if the agent correctly predicts the $i$-th value from the initial observation vector $obs$. The index of this $i$-th value is specified at the last step observation vector in $obs$[1]. To succeed, the agent must remember the sequence of observations and use this information to make an accurate prediction at the final step.
F.4 Minigrid-Memory
@@
Minigrid-Memory (Chevalier-Boisvert et al., 2023) is a two-dimensional grid-based environment that features a T-shaped maze with a small room at the beginning of the corridor, containing an object. The agent starts at a random position within the corridor. Its task is to reach the room, observe and memorize the object, then proceed to the junction at the maze’s end and turn towards the direction where an identical object is located. The reward function is defined as $R_{t}=1-0.9\times\frac{t}{T}$ for a successful attempt; otherwise, the agent receives zero reward. The episode terminates when the agent makes a choice at the junction or exceeds a time limit of steps.
F.5 Ballet
@@
In the Ballet environment (Lampinen et al., 2021) tasks take place in an $11\times 1$ 11 tiled room, consisting of a $9\times 9$ central area surrounded by a one-tile-wide wall. Each tile is upsampled to 9 pixels, resulting in a $99\times 9$ 99 pixel input image. The agent is initially placed at the center of the room, while dancers are randomly positioned in one of 8 possible locations around it. Each dancer has a distinct shape and color, selected from 15 possible shapes and 19 colors, ensuring uniqueness. These visual features serve only for identification and do not influence behavior. The agent itself is always represented as a white square. The agent receives egocentric visual observations, meaning its view is centered on its own position, which has been shown to enhance generalization.
F.6 Passive T-Maze
@@
The Passive T-Maze environment (Ni et al., 2023) consists of a corridor leading to a junction that connects two possible goal states. The agent starts at a designated position and can move in four directions: left, right, up, or down. At the beginning of each episode, one of the two goal states is randomly assigned as the correct destination. The agent observes this goal location before starting its movement. The agent stays in place if it attempts to move into a wall. To succeed, the agent must navigate to the correct goal based on its initial observation. The optimal strategy involves moving through the corridor towards the junction and then selecting the correct path.
F.7 ViZDoom-Two-Colors
@@
The ViZDoom-Two-Colors (Sorokin et al., 2022) is an environment where an agent is placed in a room with constantly depleting health. The room contains red and green objects, one of which restores health (+1 reward), while the other reduces it (-1 reward). The beneficial color is randomly assigned at the beginning of each episode and indicated by a column. The environment is structured so that the agent must memorize the column’s color to collect the correct items. Initially, the column remains visible, but in a harder variant, it disappears after 45 steps, increasing the memory requirement. To succeed, the agent must maximize survival by collecting beneficial objects while avoiding harmful ones.
F.8 Memory Maze
@@
The Memory Maze environment Pasukonis et al. (2022) is a procedurally generated 3D maze. Each episode, the agent spawns in a new maze with multiple colored objects placed in fixed locations. The agent receives a first-person view and a prompt indicating the color of the target object. It must navigate the maze, memorize object positions, and return to them efficiently. The agent receives a reward of 1 for reaching the correct object and no reward for incorrect objects.
F.9 MemoryGym Mortar Mayhem
@@
Mortar Mayhem (Pleines et al., 2023) is a grid-based environment where the agent must memorize and execute a sequence of commands in the correct order. The environment consists of a finite grid, where the agent initially observes a series of movement instructions and then attempts to reproduce them accurately. Commands include movements to adjacent tiles or remaining in place. The challenge lies in the agent’s ability to recall and execute a growing sequence of instructions, with failure resulting in episode termination. A reward of +0.1 is given for each correctly executed command
F.10 MemoryGym Mystery Path

### v2

@@ Appendix A MIKASA-Robo Implementation Details
An example of running the environment from the MIKASA-Robo benchmark is shown in 1. For ease of debugging, we also added various wrappers (found in mikasa_robo_suite/utils/wrappers/) that display useful information about the episode on the video (2). Thus, RenderStepInfoWrapper() displays the current step in the environment; DebugRewardWrapper() displays information about the full reward at the current step in the environment; DebugRewardWrapper() displays information about each component that generates the reward function at the current step. In addition, we also added task-specific wrappers for each environment. For example, RememberColorInfoWrapper() displays the target color of the cube in the RememberColor-v0 task, and ShellGameRenderCupInfoWrapper() displays which mug the ball is actually under in the ShellGame-v0 task.
@@
# pip install mikasa_robo_suite
import mikasa_robo_suite
from mikasa_robo_suite.utils.wrappers import StateOnlyTensorToDictWrapper
from tqdm.notebook import tqdm
import torch
import gymnasium as gym
# Create the environment via gym.make()
# obs_mode="rgb" for modes "RGB", "RGB+joint", "RGB+oracle" etc.
# obs_mode="state" for mode "state"
episode_timeout = 90
env = gym.make("RememberColor9-v0", num_envs=512 obs_mode="rgb", render_mode="all")
env = StateOnlyTensorToDictWrapper(env) # * always use this wrapper!
obs, _ = env.reset(seed=42)
print(obs.keys())
for i in tqdm(range(episode_timeout)):
action = torch.from_numpy(env.action_space.sample())
obs, reward, terminated, truncated, info = env.step(action)
env.close()
@@
import mikasa_robo_suite, torch
from mikasa_robo_suite.dataset_collectors.get_mikasa_robo_datasets import env_info
import gymnasium as gym
from mani_skill.utils.wrappers import RecordEpisode
from IPython.display import Video
env = gym.make("RememberColor9-v0", num_envs=512, obs_mode="rgb", render_mode="all")
state_wrappers_list, episode_timeout = env_info("RememberColor9-v0")
for wrapper_class, wrapper_kwargs in state_wrappers_list:
env = wrapper_class(env, **wrapper_kwargs)
env = RecordEpisode(env, f"./videos", max_steps_per_video=episode_timeout)
obs, _ = env.reset(seed=42)
for i in range(episode_timeout):
action = torch.from_numpy(env.action_space.sample())
obs, reward, terminated, truncated, info = env.step(action)
Video(f"./videos/0.mp4", embed=True, width=640)
env.close()
Appendix B MIKASA-Robo Datasets for Offline RL
To train Offline RL baselines on camera images (in “RGB” mode) with sparse rewards (success condition), we collected datasets for each of the 32 MIKASA-Robo tasks. Datasets were collected using a PPO-MLP agent trained to SR=100% in “state” mode (i.e., with full information about the task being solved) with sparse rewards (success condition). Thus, each dataset is represented by 1000 successful trajectories, where each trajectory consists of:
  1. 1.
    “rgb” (shape: $(T,128,128,6)$) - two RGB images (view from above and from the gripper)
  2. 2.
    “joints” (shape: $(T,25)$) - Tool Center Point (TCP) position and rotation, and joint positions and velocities
  3. 3.
    “action” (shape: $(T,8)$) - action (8-dimensional vector)
  4. 4.
    “reward” (shape: $(T,)$) - (dense) reward for each step
  5. 5.
    “success” (shape: ($T,$)) - (sparse) success flag for each step
  6. 6.
    “done” (shape: ($T,$)) - done flag for each step
These datasets are available for download from the project website. We have also published the weights of the PPO-MLP agent used to collect the dataset, as well as scripts for collecting datasets of any size, to our repository.
Appendix C MIKASA-Base Implementation Details
@@
import mikasa_base
import gymnasium as gym
# use pre-defined env
# env_id = "MemoryLengthEasy-v0"
# env_kwargs = None
# create env using custom parameters
env_id = "MemoryLength-v0"
env_kwargs = {"memory_length": 10, "num_bits": 1}
seed = 123
env = gym.make(env_id, env_kwargs)
obs, _ = env.reset(seed=seed)
for i in range(11):
action = env.action_space.sample()
next_obs, reward, terminations, truncations, infos = env.step(action)
env.close()
Appendix D MIKASA-Robo setup for VLA baselines
For experiments involving Vision-Language-Action (VLA) models, we focused on a representative subset of spatial and object memory tasks from MIKASA-Robo. For each task, we generated a dataset of 250 episodes using an oracle PPO policy with full access to the environment state. At every timestep, the policy recorded two synchronized RGB frames (one from the static “base” camera and one from the robot’s wrist camera) along with the corresponding end-effector control actions ( pd_ee_delta_pose controller from [87]). Each task was also paired with a concise language instruction (see Table 5).
All VLA baselines were trained for 50000 iterations and evaluated independently on each task. Complete training/evaluation scripts, language instruction templates, and detailed model hyperparameter settings are provided in the accompanying supplementary code.
Table 5: Tasks configurations for fine-tuning VLA models. The table lists the task ID, number of evaluation steps (T), and the associated language instruction
Task T Language instruction RememberColor3/5/9-v0 60 Remember the color of the cube and then pick the matching one ShellGameTouch-v0 90 Memorize the position of the cup covering the ball, then pick that cup InterceptMedium-v0 90 Track the ball’s movement, estimate its velocity, then aim the ball at the target
Appendix E Memory Mechanisms in RL
@@
There are several approaches to incorporating memory into RL, including recurrent neural networks (RNNs) [76, 37, 14] which uses hidden states to store information from previous steps [93, 34], state-space models (SSMs) [28, 83, 27] which uses system state to store historical information [31, 77], transformers [92] which uses attention mechanism to capture sequential dependencies inside the context window [72, 54, 68], graph neural networks (GNNs) [98] which uses graphs to store information [99, 44] etc. Popular agents with memory mechanisms are summarized in Table 2.
Appendix F Classic baselines performance on the MIKASA-Robo benchmark
@@
Appendix G Experiments Reproducing and Compute Resources
All baselines were trained and evaluated under a reproducible standardized setup. Training for every algorithm was performed on a single NVIDIA A100 GPU. For evaluation, each task was run for 100 independent episodes, with environment and agent random seeds ranging from 1 to 100. We report performance metrics as the mean success rate $\pm$ the standard error of the mean (SEM) over these 100 trials.
Appendix H MIKASA-Robo Detailed Tasks Description
@@
Table 6: Results for Offline RL baselines. The table shows comparison of transformer-based baselines (RATE, DT), behavior cloning (BC), classic Offline RL baselines (CQL), and Diffusion Policy (DP) on all 32 tasks from the MIKASA-Robo benchmark. Results are presented as mean ± sem across the three runs, where each run is averaged over 100 episodes and sem is the standard error of the mean. Training was performed using only RGB observations (two cameras: top view and gripper view) and using sparse rewards (success once condition). The results show that even models with memory (RATE, DT) are not able to solve most of the benchmark problems, which makes it challenging and promising for further validation of the algorithm.
# Environment RATE DT BC CQL DP
1 ShellGameTouch-v0 0.92±0.01 0.53±0.07 0.28±0.01 0.16±0.04 0.18±0.02
2 ShellGamePush-v0 0.78±0.06 0.62±0.14 0.27±0.01 0.25±0.01 0.22±0.03
3 ShellGamePick-v0 0.02±0.01 0.00±0.00 0.01±0.01 0.00±0.00 0.01±0.00
4 InterceptSlow-v0 0.23±0.02 0.40±0.02 0.37±0.06 0.25±0.01 0.33±0.05
5 InterceptMedium-v0 0.32±0.02 0.56±0.01 0.31±0.14 0.03±0.01 0.68±0.02
6 InterceptFast-v0 0.30±0.04 0.36±0.04 0.03±0.02 0.02±0.02 0.21±0.05
7 InterceptGrabSlow-v0 0.09±0.03 0.00±0.00 0.28±0.18 0.03±0.00 0.03±0.01
8 InterceptGrabMedium-v0 0.09±0.03 0.00±0.00 0.11±0.02 0.08±0.04 0.03±0.01
9 InterceptGrabFast-v0 0.14±0.03 0.11±0.03 0.09±0.02 0.08±0.03 0.18±0.02
10 RotateLenientPos-v0 0.11±0.04 0.01±0.01 0.15±0.03 0.16±0.02 0.11±0.02
11 RotateLenientPosNeg-v0 0.29±0.03 0.05±0.02 0.22±0.01 0.12±0.02 0.14±0.05
12 RotateStrictPos-v0 0.03±0.02 0.05±0.04 0.01±0.00 0.03±0.01 0.06±0.02
13 RotateStrictPosNeg-v0 0.08±0.01 0.05±0.03 0.04±0.02 0.04±0.02 0.15±0.01
14 TakeItBack-v0 0.42±0.24 0.08±0.04 0.33±0.10 0.04±0.01 0.05±0.02
15 RememberColor3-v0 0.65±0.04 0.01±0.01 0.27±0.03 0.29±0.01 0.32±0.01
16 RememberColor5-v0 0.13±0.03 0.07±0.05 0.12±0.01 0.15±0.02 0.10±0.02
17 RememberColor9-v0 0.09±0.02 0.01±0.01 0.12±0.02 0.15±0.01 0.17±0.01
18 RememberShape3-v0 0.21±0.04 0.05±0.04 0.31±0.04 0.20±0.10 0.32±0.05
19 RememberShape5-v0 0.17±0.04 0.04±0.04 0.18±0.01 0.15±0.00 0.21±0.04
20 RememberShape9-v0 0.05±0.00 0.05±0.02 0.10±0.02 0.14±0.01 0.11±0.02
21 RememberShapeAndColor3x2-v0 0.14±0.02 0.04±0.02 0.13±0.02 0.11±0.05 0.14±0.02
22 RememberShapeAndColor3x3-v0 0.08±0.03 0.06±0.06 0.09±0.02 0.09±0.02 0.16±0.01
23 RememberShapeAndColor5x3-v0 0.07±0.02 0.01±0.01 0.09±0.01 0.09±0.02 0.11±0.03
24 BunchOfColors3-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
25 BunchOfColors5-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
26 BunchOfColors7-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
27 SeqOfColors3-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
28 SeqOfColors5-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
29 SeqOfColors7-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
30 ChainOfColors3-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
31 ChainOfColors5-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
32 ChainOfColors7-v0 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00
@@ H.12 ChainOfColors-v0
Appendix I MIKASA-Base Benchmark Tasks Description
@@ I.1 Memory Cards
The Memory Cards environment [19] is a memory game environment with 5 randomly shuffled pairs of hidden cards. At each step, the agent sees one revealed card and must find its matching pair. A correct guess removes both cards; otherwise, the card is hidden again, and a new one is revealed. The game continues until all pairs are removed.
I.2 Numpad
@@
The Numpad environment [38] consists of an $N\times N$ grid of tiles. The agent controls a ball that rolls between tiles. At the beginning of an episode, a random sequence of $n$ neighboring tiles (excluding diagonals) is selected, and the agent must follow this sequence in the correct order. The environment is structured so that pressing the correct tile lights it up, while pressing an incorrect tile resets progress. A reward of +1 is given for the first press of each correct tile after a reset. The episode ends after a fixed number of steps. To succeed, the agent must memorize the sequence and navigate it correctly without mistakes. The ability to “jump” over tiles is not available.
I.3 BSuite MemoryLength
@@
The MemoryLength environment [70] represents a sequence of observations, where at each step, the observation takes a value of either +1 or -1. The environment is structured so that a reward is given only at the final step if the agent correctly predicts the $i$-th value from the initial observation vector $obs$. The index of this $i$-th value is specified at the last step observation vector in $obs$[1]. To succeed, the agent must remember the sequence of observations and use this information to make an accurate prediction at the final step.
I.4 Minigrid-Memory
@@
Minigrid-Memory [12] is a two-dimensional grid-based environment that features a T-shaped maze with a small room at the beginning of the corridor, containing an object. The agent starts at a random position within the corridor. Its task is to reach the room, observe and memorize the object, then proceed to the junction at the maze’s end and turn towards the direction where an identical object is located. The reward function is defined as $R_{t}=1-0.9\times\frac{t}{T}$ for a successful attempt; otherwise, the agent receives zero reward. The episode terminates when the agent makes a choice at the junction or exceeds a time limit of steps.
I.5 Ballet
@@
In the Ballet environment [54] tasks take place in an $11\times 1$ 11 tiled room, consisting of a $9\times 9$ central area surrounded by a one-tile-wide wall. Each tile is upsampled to 9 pixels, resulting in a $99\times 9$ 99 pixel input image. The agent is initially placed at the center of the room, while dancers are randomly positioned in one of 8 possible locations around it. Each dancer has a distinct shape and color, selected from 15 possible shapes and 19 colors, ensuring uniqueness. These visual features serve only for identification and do not influence behavior. The agent itself is always represented as a white square. The agent receives egocentric visual observations, meaning its view is centered on its own position, which has been shown to enhance generalization.
I.6 Passive T-Maze
@@
The Passive T-Maze environment [68] consists of a corridor leading to a junction that connects two possible goal states. The agent starts at a designated position and can move in four directions: left, right, up, or down. At the beginning of each episode, one of the two goal states is randomly assigned as the correct destination. The agent observes this goal location before starting its movement. The agent stays in place if it attempts to move into a wall. To succeed, the agent must navigate to the correct goal based on its initial observation. The optimal strategy involves moving through the corridor towards the junction and then selecting the correct path.
I.7 ViZDoom-Two-Colors
@@
The ViZDoom-Two-Colors [84] is an environment where an agent is placed in a room with constantly depleting health. The room contains red and green objects, one of which restores health (+1 reward), while the other reduces it (-1 reward). The beneficial color is randomly assigned at the beginning of each episode and indicated by a column. The environment is structured so that the agent must memorize the column’s color to collect the correct items. Initially, the column remains visible, but in a harder variant, it disappears after 45 steps, increasing the memory requirement. To succeed, the agent must maximize survival by collecting beneficial objects while avoiding harmful ones.
I.8 Memory Maze
@@
The Memory Maze environment [73] is a procedurally generated 3D maze. Each episode, the agent spawns in a new maze with multiple colored objects placed in fixed locations. The agent receives a first-person view and a prompt indicating the color of the target object. It must navigate the maze, memorize object positions, and return to them efficiently. The agent receives a reward of 1 for reaching the correct object and no reward for incorrect objects.
I.9 MemoryGym Mortar Mayhem
@@
Mortar Mayhem [75] is a grid-based environment where the agent must memorize and execute a sequence of commands in the correct order. The environment consists of a finite grid, where the agent initially observes a series of movement instructions and then attempts to reproduce them accurately. Commands include movements to adjacent tiles or remaining in place. The challenge lies in the agent’s ability to recall and execute a growing sequence of instructions, with failure resulting in episode termination. A reward of +0.1 is given for each correctly executed command
I.10 MemoryGym Mystery Path